6. In the town of Troy, NY there are 51,340 residents. 53% of the residents are male. A new illness never before described emerged in the town during 2018. It was norovirus-like but had an added symptom of a full body rash. Between January-March, 345 women and 125 men came down with the illness. From April to June, 45 men and 127 women contracted the illness. July-September, 23 men and 157 women were sick with this new illness. October - December, Troy residents saw a decrease in the overall cases, 5 men and 25 women were diagnosed with the disease. Overall, 30 people died due to this illness. Looking at the health data we know that there were 1,750 total deaths due to other causes in the town of Troy during 2018.

a) Calculate the incidence rate of this new illness during 2018 per 10,000 population in Troy, NY.
b) Calculate the incidence rates per 10,000 population for each quarterly period.
c) Calculate the sex specific incidence rates per 10,000 population during 2018.
d) Determine the ratio of female cases to male cases.
e) Calculate prevalence of this new illness.
f) Calculate point prevalence on June 30.
g) What is the case fatality rate of this new illness?

it a case study all the info u need is there
Answer #1

6. The incidence rate is the number of new cases per population in a given time period.

a). The incidence rate per 10,000 population is 852/ 51340 X 10,000 = 165.9

b). Incidence rates pr 10,000 Quaterly are:

Jan. - March = 470 / 51340 X 10,000 = 91.5

April - June = 172 / 50870 X 10,000 = 33.8

July - Sep. = 180 / 50698 X 10,000 = 35.5

Oct. - Dec. = 30 / 50518 X 10,000 = 5.9

c). Sex specific incidence rates per 10,000 population;

Men = 198 / 51340 X 10,000 = 38.5

Women = 654 / 51340 X 10,000 = 127.4

d). The ratio of female cases to male cases = 654 / 198 = 3.3

e). Prevalence of this new illness is = 852 / 51340 = 0.0166

f). Point prevalence on june 30 = 642 / 51340 = 0.0125    = 1.25%

g). The case fatality rate of new illness = 30 / 852 = 0.035 = 3.5%